Calculate the enthalpy change associated with the following process:

a) m-cresol is heated from 25°C to 100°C

b) Ethylene glycol is cooled from 15° to 120°C

c) Succinic acid is heated from 15°C to 65°C

d) Air is cooled from 15°C to 65°C
Answer #1

Enthalpy change expression will be: Ah=C,AT Ah=C,(Iz-I) As the mass or moles of the compounds are not given in the question s(b) Ethylene glycol is cooled from 15°C to 120°C. The heat capacity of Ethylene glycol in between the given temperature range(c) Succinic acid is heated from 15°C to 65°C. The heat capacity of Succinic acid is, C, =(0.248+0.00153T), and this value of(d) Air is cooled from 15°C to 65°C. Ah=C,d1 =(a+bT+cT? +dT?) dT = f(a+b7 +cT2 +dT°)dt 28.94 +0.4147x10-2T +0.3191x10-12-1.96